# When Hyperspectral Image Classification Meets Diffusion Models: An Unsupervised Feature Learning Framework

### v1

@@ 16 Stone-Steel-Towers 9 84
TABLE II: Land-cover types, the number of labeled training samples and testing samples of the Pavia University dataset.
@@ 20 Stadium seats 341 6483
TABLE IV: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$ as well as the accuracies for each class on the Indian Pines dataset. The best one is shown in bold.
Class CNN Based Methods Transformer Based Methods Unsupervised Methods Diff-HSI
@@
2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE UMSDFL
1 65.85 58.54 94.14 63.00 95.12 97.56 72.97 99.10 99.51
2 99.77 76.19 97.84 92.35 97.67 98.05 88.50 96.10 98.89
3 81.66 77.64 97.54 86.86 98.87 98.66 87.20 95.39 99.84
4 96.71 52.11 90.70 88.96 91.55 95.31 84.90 97.24 99.81
5 85.75 93.56 97.75 92.49 96.32 95.17 90.28 94.12 98.85
6 97.87 98.17 99.24 99.12 99.54 99.85 97.97 99.25 100.00
7 100.00 36.00 81.60 52.50 100.00 100.00 56.52 88.46 98.40
8 100.00 98.60 100.00 99.16 100.00 100.00 99.48 100.00 100.00
9 50.00 55.56 74.44 41.18 88.89 100.00 87.50 94.44 100.00
10 35.54 82.86 94.77 93.16 97.71 94.29 86.80 95.84 98.88
11 88.01 90.45 98.87 92.27 98.69 99.37 96.68 99.29 99.63
12 98.13 62.55 97.83 85.44 98.13 96.63 80.83 93.37 98.91
13 99.46 88.65 99.24 99.02 97.28 100.00 100.00 100.00 99.89
14 99.91 99.39 99.18 96.73 99.91 97.89 99.90 95.25 100.00
15 91.35 86.17 93.95 83.41 98.84 97.12 96.80 99.15 99.25
16 86.90 45.24 98.33 93.50 95.54 94.05 84.00 100.00 98.10
OA (%) 87.77 85.42 97.75 92.31 97.47 97.95 92.69 97.02 99.46
AA (%) 86.06 75.10 94.71 84.95 96.57 97.75 88.15 96.00 99.37
$\kappa$ 0.8603 0.8324 0.9743 0.9124 0.9711 0.9766 0.9162 0.9660 0.9938
TABLE V: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$, as well as the accuracies for each class on the Pavia University dataset. The best one is shown in bold.
Class CNN Based Methods Transformer Based Methods Unsupervised Methods Diff-HSI
2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE UMSDFL
1 99.68 97.02 98.81 96.21 99.33 99.38 94.20 99.62 100.00
2 99.41 99.97 99.83 99.64 99.92 99.80 99.58 99.98 99.99
3 86.56 92.98 92.45 87.65 98.29 98.35 78.83 91.02 100.00
4 98.18 97.53 98.32 96.64 98.49 99.52 97.53 98.40 99.51
5 99.84 99.06 99.65 99.97 99.53 100.00 100.00 100.00 100.00
6 100.00 99.10 99.43 99.56 100.00 99.75 95.30 98.32 100.00
7 99.84 79.10 99.76 90.30 99.13 99.60 97.42 99.61 100.00
8 100.00 97.34 99.32 94.60 98.05 98.63 96.87 98.36 99.82
9 98.22 95.22 99.82 98.49 95.44 99.33 98.61 99.89 99.96
OA (%) 98.86 97.88 99.10 97.54 99.21 99.53 96.77 99.02 99.94
AA (%) 97.97 95.26 98.60 95.88 98.69 99.37 95.37 98.36 99.92
$\kappa$ 0.9848 0.9719 0.9881 0.9674 0.9915 0.9937 0.9571 0.9870 0.9993
@@ 2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE UMSDFL
TABLE VI: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$ as well as the accuracies for each class on the Houston2018 dataset. The best one is shown in bold.
Class CNN Based Methods Transformer Based Methods Unsupervised Methods Diff-HSI
@@
2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE UMSDFL
1 84.33 82.02 86.30 92.36 79.93 79.50 91.60 88.99 89.41
2 95.98 96.64 95.32 95.08 93.44 96.55 93.92 97.53 96.24
3 96.92 96.00 99.72 96.21 99.66 100.00 97.60 99.23 99.88
4 98.13 95.55 97.49 98.48 96.64 97.62 95.92 97.98 99.26
5 91.26 79.16 86.09 91.87 90.11 95.01 85.79 92.51 96.86
6 99.91 97.51 98.48 99.62 99.62 99.91 98.50 99.58 100.00
7 80.63 71.94 93.68 27.01 85.45 95.65 61.15 96.85 97.63
8 94.29 88.62 91.62 95.42 98.72 99.18 91.15 94.20 99.62
9 99.54 92.80 97.88 98.60 99.09 99.35 95.12 99.07 99.63
10 95.44 71.61 81.72 88.22 91.15 92.64 76.72 88.61 95.89
11 77.86 73.17 69.44 27.01 80.97 85.73 70.22 82.28 93.02
12 15.26 11.10 0.51 31.46 41.69 33.43 4.79 42.76 50.23
13 87.30 69.12 84.59 91.79 94.49 96.79 90.52 93.26 97.72
14 94.97 96.18 89.65 92.91 96.97 99.17 87.23 97.02 99.23
15 95.40 98.98 99.34 99.33 99.30 99.92 99.06 99.59 99.81
16 93.62 90.40 91.00 96.36 97.84 98.15 92.01 96.36 99.69
17 86.33 20.86 0.00 22.78 69.21 90.65 0.00 100.00 94.10
18 96.64 89.05 93.66 91.61 93.07 97.85 90.43 94.24 99.46
19 99.24 95.45 96.92 96.53 97.97 99.84 96.09 99.59 99.89
20 92.77 93.92 99.12 99.77 99.96 100.00 97.27 99.89 99.94
OA (%) 94.84 86.88 91.61 90.65 95.48 96.69 90.34 95.38 98.07
AA (%) 88.79 80.50 82.63 80.75 90.26 92.85 80.76 92.98 95.37
$\kappa$ 0.9325 0.8313 0.8906 0.8784 0.9412 0.9570 0.8751 0.9398 0.9748
@@ IV-B2 Comparison with State-of-the-art Methods
To demonstrate the effectiveness of our proposed method, we compare our classification performance with several state-of-the-art approaches, including CNN-based and transformer-based supervised methods, and unsupervised methods. We use the most effective setting for each of these methods.
  • The 2-D CNN architecture contains three 2-D convolution blocks and a softmax layer. Each convolution block consists of a 2-D convolution layer, a BN layer, an avg-pooling layer, and a ReLU activation function.
  • The 3-D CNN contains three 3-D convolution blocks and a softmax layer. Each 3-D convolution block consists of a 3-D convolution layer, a BN layer, a ReLU activation function, and a 3-D convolution layer with step size 2.
@@
  • For SF[16], there are four encoder blocks in the ViT-based network. In the transformer framework of SF, group-wise spectral embedding and cross-layer adaptive fusion modules are adopted to capture local spectral representations from neighboring bands.
@@
  • The 3DCAE[21] is an unsupervised method using the encoder-decoder backbone with 3D convolution operation to learn spectral-spatial features.
  • The UMSDFL[22] is an unsupervised method using encoder and decoder with convolutional layers to learn spectral-spatial features. A clustering branch and a multi-layer fusion module is designed to enhance the features.
@@ IV-B3 Implementation Details
The proposed Diff-HSI was implemented using the Pytorch framework. The patch size is set to $48\times 48$, and the dimension of PCA is set to 10. In the diffusion-pretraining procedure, we use Kullback-Leibler Divergence Loss as the loss function. And the Adam optimizer is adopted with a batch size of 128 and a learning rate of 1e-4. In the second stage, the cross-entropy loss is used in lightweight classifiers. $m$, $n$, and $k$ are set to be 20, 3, and 5, respectively. We adopt the Adam Optimizer and the Cosine Annealing as our training schedule. The original learning rate and minimum learning rate are set to be 1e-4 and 5e-6, respectively. The number of epochs is set to 100 for all datasets. Fairly, we calculate the results by averaging the results of ten repeated experiments with different training sample selections.
@@ IV-C Quantitative Results and Analysis
Quantitative classification results in terms of class-specific accuracy, OA, AA, and $\kappa$ of the compared methods on the Indian Pines, PaviaU, and Houston 2018 datasets are listed in Table IV, V, and VI, respectively. And the classification maps of all methods are shown in Fig. 4, 5 and 6.
@@
Compared with other methods, our proposed Diff-HSI achieves the highest OA, AA, and $\kappa$ on the three datasets. According to the results, the CNN-based supervised methods obtain good performance owing to the ability of capturing local spatial information. Besides, since transformers are capable of capturing sequential information, transformer-based supervised methods also achieve competitive performance. The GAHT method achieves the best results in CNN-based and transformer-based supervised methods because it balances the local information and the global dependencies. However, the methods that learn explicit mappings have bottlenecks in learning label-agnostic information and complex relations, which limits performance. Some unsupervised methods are proposed to tackle the problem by learning representative features through an encoder-decoder network. Limited by the model architecture, the features learned are not discriminative and rich enough for HSI classification lacking high-level information. Therefore, the performance of these methods is even lower than that of some explicit learning methods. Our proposed Diff-HSI introduces the diffusion model to HSI classification, which has the powerful ability to model complex relations. Furthermore, the features learned by our method are of great diversity, enriching the representation capability. Thus, our Diff-HSI outperforms all the previous methods on the three datasets. Notably, the classification performance of the Houston2018 Dataset is largely improved compared with the previous SOTA method in terms of OA (98.07% versus 96.69%), AA (95.37% versus 92.98%), and $\kappa$ (0.9748 versus 0.9570), which especially demonstrate our effectiveness.

### v2

@@ 16 Stone-Steel-Towers 9 84
TABLE II: Land-cover types, the number of labeled training samples and testing samples of the PaviaU dataset.
@@ 9 Shadows 47 900
After obtaining the multi-timestep multi-stage selective representation, a lightweight network is needed to predict the classification label. Inspired by [46], we train an ensemble of lightweight linear classifiers that takes the dynamic pixel representations as inputs and predicts the classification label of each pixel. Specifically, each classifier is trained independently, consisting of two hidden layers with ReLU activation and batch normalization. When testing a sample, the final predicted label is obtained by majority voting of the ensemble of pixel classifiers, as illustrated in Fig. 2. This method brings more stability of prediction with a very small cost since the parameters of each classifier are very limited.
IV Experiments and results
In this section, we first describe four well-known HSI datasets, including the Indian Pines dataset, the Pavia University dataset, the Houston 2018 dataset, and the WHU-Hi-Longkou dataset. The experimental setting is then introduced including evaluation metrics, a brief introduction of compared state-of-art methods, and implementation details. Then, we conduct quantitative experiments and ablation analysis to evaluate our proposed method.
IV-A Datasets Description
IV-A1 Indian Pines
IV-A2 Pavia University
The Pavia University (PaviaU) dataset was collected in 2003 by the reflective optics system imaging spectrometer (ROSIS-3) sensor over a part of the city of Pavia, Italy. The dataset consists of $610\times 340$ pixels with a spatial resolution of 1.3 m and 115 spectral bands in the wavelength range of 430 to 860 nm. 103 out of 115 bands are used for classification after removing 12 noisy bands. The image contains a large number of background pixels, and only 42776 labeled pixels are divided into 9 classes, including asphalt, meadows, gravel, and so on. We use 5% of the labeled samples for training and the rest for testing. The class name and the number of training and testing samples are listed in Table II.
IV-A3 Houston 2018
The Houston 2018 dataset, identified as the 2018 IEEE GRSS DFC dataset, was gathered in 2018 by the National Center for Airborne Laser Mapping (NCALM) over the University of Houston campus and its neighboring urban area, including HSI, multispectral LiDAR, and very high-resolution RGB images. The HSI dataset consists of $601\times 2384$ pixels with a spatial resolution of 1 m and 48 spectral bands in the wavelength range of 380 to 1050 nm. It contains 504856 labeled pixels and 20 classes of interest. We use 5% of the labeled samples for training and the rest for testing. The class name and the number of training and testing samples are listed in Table III.
IV-A4 WHU-Hi-Longkou
The WHU-Hi-Longkou dataset was acquired in 2018 by an 8-mm focal length Headwall Nano-Hyperspec imaging sensor equipped on a DJ-innovations Matrice 600 Pro UAV platform. It consists of $550\times 400$ pixels with a spatial resolution of 0.463 m and 270 spectral bands in the wavelength range of 400 to 1000 nm. It contains 204542 labeled samples and 9 object classes. We use 0.5% of the labeled samples for training and the rest for testing. The class name and the number of training and testing samples are listed in Table IV.
@@ 20 Stadium seats 341 6483
TABLE IV: Land-cover types, the number of labeled training samples and testing samples of the WHU-Hi-Longkou dataset.
Class Land Cover Type Training Testing
1 Corn 172 34339
2 Cotton 42 8332
3 Sesame 15 3016
4 Broad-leaf soybean 316 62896
5 Narrow-leaf soybean 21 4130
6 Rice 59 11795
7 Water 335 66721
8 Roads and houses 36 7088
9 Mixed weed 26 5203
Total 1022 203520
@@ 9 Mixed weed 26 5203
TABLE V: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$ as well as the accuracies for each class on the Indian Pines dataset. The best results are shown in bold.
@@
Class 2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE 3DAES UMSDFL SpectralDiff MTMSD
1 65.85 58.54 94.14 63.00 95.12 97.56 72.97 100.00 99.10 100.00 100.00
2 99.77 76.19 97.84 92.35 97.67 98.05 88.50 89.34 96.10 97.90 99.52
3 81.66 77.64 97.54 86.86 98.87 98.66 87.20 95.98 95.39 98.93 99.01
4 96.71 52.11 90.70 88.96 91.55 95.31 84.90 95.31 97.24 100.00 99.62
5 85.75 93.56 97.75 92.49 96.32 95.17 90.28 88.74 94.12 94.02 98.62
6 97.87 98.17 99.24 99.12 99.54 99.85 97.97 99.09 99.25 99.54 99.97
7 100.00 36.00 81.60 52.50 100.00 100.00 56.52 56.00 88.46 100.00 99.20
8 100.00 98.60 100.00 99.16 100.00 100.00 99.48 99.77 100.00 100.00 100.00
9 50.00 55.56 74.44 41.18 88.89 100.00 87.50 100.00 94.44 100.00 100.00
10 35.54 82.86 94.77 93.16 97.71 94.29 86.80 87.77 95.84 98.51 98.79
11 88.01 90.45 98.87 92.27 98.69 99.37 96.68 96.88 99.29 99.77 99.67
12 98.13 62.55 97.83 85.44 98.13 96.63 80.83 83.71 93.37 91.76 98.84
13 99.46 88.65 99.24 99.02 97.28 100.00 100.00 100.00 100.00 99.46 99.78
14 99.91 99.39 99.18 96.73 99.91 97.89 99.90 99.30 95.25 99.91 99.86
15 91.35 86.17 93.95 83.41 98.84 97.12 96.80 98.56 99.15 98.27 99.42
16 86.90 45.24 98.33 93.50 95.54 94.05 84.00 97.62 100.00 98.81 98.10
OA (%) 87.77 85.42 97.75 92.31 97.47 97.95 92.69 94.34 97.02 98.54 99.45
AA (%) 86.06 75.10 94.71 84.95 96.57 97.75 88.15 93.00 96.00 98.56 99.40
$\kappa$ 0.8603 0.8324 0.9743 0.9124 0.9711 0.9766 0.9162 0.9353 0.9660 0.9833 0.9937
Figure 5: Classification maps obtained by different methods on the Indian Pines dataset. (a) Ground truth. (b) 2-D CNN (OA=87.77%). (c) 3-D CNN (OA=85.42%). (d) SSRN (OA=97.75%). (e) SF (OA=92.31%). (f) SSFTT (OA=97.47%). (g) GAHT (OA=97.95%). (h) 3DCAE (OA=92.69%). (i) 3DAES (OA=94.34%). (j) UMSDFL (OA=97.02%). (k) SpectralDiff (OA=98.54%). (l) MTMSD (OA=99.45%).
@@
TABLE VI: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$, as well as the accuracies for each class on the PaviaU dataset. the best results are shown in bold.
@@
Class 2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE 3DAES UMSDFL SpectralDiff MTMSD
1 99.68 97.02 98.81 96.21 99.33 99.38 94.20 97.94 99.62 99.89 100.00
2 99.41 99.97 99.83 99.64 99.92 99.80 99.58 99.00 99.98 100.00 100.00
3 86.56 92.98 92.45 87.65 98.29 98.35 78.83 89.57 91.02 98.75 100.00
4 98.18 97.53 98.32 96.64 98.49 99.52 97.53 98.63 98.40 97.35 99.59
5 99.84 99.06 99.65 99.97 99.53 100.00 100.00 100.00 100.00 95.93 100.00
6 100.00 99.10 99.43 99.56 100.00 99.75 95.30 97.11 98.32 100.00 100.00
7 99.84 79.10 99.76 90.30 99.13 99.60 97.42 94.77 99.61 100.00 100.00
8 100.00 97.34 99.32 94.60 98.05 98.63 96.87 98.20 98.36 99.77 99.87
9 98.22 95.22 99.82 98.49 95.44 99.33 98.61 97.44 99.89 94.44 99.79
OA (%) 98.86 97.88 99.10 97.54 99.21 99.53 96.77 97.92 99.02 99.46 99.95
AA (%) 97.97 95.26 98.60 95.88 98.69 99.37 95.37 96.96 98.36 98.46 99.92
$\kappa$ 0.9848 0.9719 0.9881 0.9674 0.9915 0.9937 0.9571 0.9725 0.9870 0.9929 0.9994
Figure 6: Classification maps obtained by different methods on the PaviaU dataset. (a) Ground truth. (b) 2-D CNN (OA=98.86%). (c) 3-D CNN (OA=97.88%). (d) SSRN (OA=99.10%). (e) SF (OA=97.54%). (f) SSFTT (OA=99.21%). (g) GAHT (OA=99.53%). (h) 3DCAE (OA=96.77%). (i) 3DAES (OA=97.92%). (j) UMSDFL (OA=99.02%). (k) SpectralDiff (OA=99.46%). (l) MTMSD (OA=99.95%).
@@ IV-B2 Comparison with State-of-the-art Methods
To demonstrate the effectiveness of our proposed method, we compare our classification performance with several state-of-the-art approaches using the most effective setting for these methods.
  • The 2-D CNN[47] architecture contains three 2-D convolution blocks and a softmax layer. Each convolution block consists of a 2-D convolution layer, a BN layer, an avg-pooling layer, and a ReLU activation function.
  • The 3-D CNN[47] contains three 3-D convolution blocks and a softmax layer. Each 3-D convolution block consists of a 3-D convolution layer, a BN layer, a ReLU activation function, and a 3-D convolution layer with step size 2.
@@
  • For SF[16], group-wise spectral embedding and cross-layer adaptive fusion modules in the transformer framework are adopted to capture local spectral representations from neighboring bands.
@@
  • The 3DCAE[20] is an unsupervised method using an encoder-decoder backbone with 3D convolution operation to learn spectral-spatial features.
  • The 3DAES[48] is a semi-supervised method using an autoencoder to extract spectral-spatial features from unlabeled samples and then optimizes the siamese network and classifier using constructed sample pairs.
  • The UMSDFL[21] is an unsupervised method using encoder and decoder with convolutional layers to learn spectral-spatial features. A clustering branch and a multi-layer fusion module are designed to enhance the features.
  • The SpectralDiff[32] is a diffusion-based unsupervised method that learns diffusion features through the spectral-spatial diffusion module and feeds diffusion features into the attention-based classification module.
@@ IV-B3 Implementation Details
The proposed MTMSD was implemented using the Pytorch framework. The patch size is set to $48\times 48$, and the dimension of PCA is set to $8/N$, $N$ is the number of spectral bands of the dataset. In the diffusion-pretraining procedure, we use Kullback-Leibler Divergence Loss as the loss function. And the Adam optimizer is adopted with a batch size of 128 and a learning rate of 1e-4, training a total of 40k steps. In the feature-exploring stage, the cross-entropy loss is used in lightweight classifiers. $m$ and $K$ are set to be 19 and 5, respectively. We adopt the Adam Optimizer and the Cosine Annealing as our training schedule. The original learning rate and minimum learning rate are set to be 1e-4 and 5e-6, respectively. The number of epochs is set to 100 for all datasets. We calculate the results fairly by averaging the results of ten repeated experiments with different training sample selections.
@@ IV-C Quantitative Results and Analysis
IV-C1 Classification Results Compared with SOTA Methods
@@
Quantitative classification results in terms of class-specific accuracy, OA, AA, and $\kappa$ of the compared methods on the Indian Pines, PaviaU, Houston 2018 and Longkou datasets are listed in Table V, VI, VII and VIII, respectively. And the classification maps of all methods are shown in Fig. 5, 6, 7 and 8.
TABLE VII: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$ as well as the accuracies for each class on the Houston 2018 dataset. the best results are shown in bold.
Class 2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE 3DAES UMSDFL SpectralDiff MTMSD
1 87.12 82.02 86.30 92.36 79.93 79.50 91.60 81.80 88.99 83.47 88.87
2 92.05 96.64 95.32 95.08 93.44 96.55 93.92 93.30 97.53 92.48 96.29
3 96.92 96.00 99.72 96.21 99.66 100.00 97.60 89.23 99.23 99.38 99.93
4 98.05 95.55 97.49 98.48 96.64 97.62 95.92 94.39 97.98 97.01 99.32
5 87.25 79.16 86.09 91.87 90.11 95.01 85.79 81.95 92.51 86.25 97.01
6 95.36 97.51 98.48 99.62 99.62 99.91 98.50 97.72 99.58 99.58 100.00
7 96.44 71.94 93.68 27.01 85.45 95.65 61.15 83.40 96.85 87.35 97.40
8 97.15 88.62 91.62 95.42 98.72 99.18 91.15 91.27 94.20 98.73 99.81
9 98.24 92.80 97.88 98.60 99.09 99.35 95.12 97.12 99.07 99.31 99.72
10 93.88 71.61 81.72 88.22 91.15 92.64 76.72 76.90 88.61 87.73 96.63
11 75.85 73.17 69.44 27.01 80.97 85.73 70.22 71.14 82.28 79.33 93.96
12 12.55 11.10 0.51 31.46 41.69 33.43 4.79 3.19 42.76 37.24 55.30
13 85.24 69.12 84.59 91.79 94.49 96.79 90.52 84.62 93.26 95.64 97.87
14 77.12 96.18 89.65 92.91 96.97 99.17 87.23 95.41 97.02 98.09 99.27
15 94.45 98.98 99.34 99.33 99.30 99.92 99.06 97.31 99.59 99.47 99.92
16 93.43 90.40 91.00 96.36 97.84 98.15 92.01 87.79 96.36 98.42 99.84
17 64.75 20.86 0.00 22.78 69.21 90.65 0.00 46.04 100.00 84.17 100.00
18 91.70 89.05 93.66 91.61 93.07 97.85 90.43 83.89 94.24 93.91 99.60
19 96.88 95.45 96.92 96.53 97.97 99.84 96.09 94.55 99.59 98.47 99.99
20 99.83 93.92 99.12 99.77 99.96 100.00 97.27 96.30 99.89 100.00 100.00
OA (%) 93.38 86.88 91.61 90.65 95.48 96.69 90.34 90.39 95.38 95.28 98.29
AA (%) 86.71 80.50 82.63 80.75 90.26 92.85 80.76 82.37 92.98 90.80 96.04
$\kappa$ 0.9137 0.8313 0.8906 0.8784 0.9412 0.9570 0.8751 0.8748 0.9398 0.9385 0.9777
Figure 7: Classification maps obtained by different methods on the Houston 2018 dataset. (a) Ground truth. (b) 2-D CNN (OA=94.84%). (c) 3-D CNN (OA=86.88%). (d) SSRN (OA=91.61%). (e) SF (OA=90.65%). (f) SSFTT (OA=95.48%). (g) GAHT (OA=96.69%). (h) 3DCAE (OA=90.34%). (i) 3DAES (OA=90.39%). (j) UMSDFL (OA=95.38%). (k) SpectralDiff (OA=95.28%). (l) MTMSD (OA=98.29%).
TABLE VIII: Quantitative performance of different classification methods in terms of OA, AA, and $\kappa$, as well as the accuracies for each class on the WHU-Hi-Longkou dataset. the best results are shown in bold.
Class 2-D CNN 3-D CNN SSRN SF SSFTT GAHT 3DCAE 3DAES UMSDFL SpectralDiff MTMSD
1 99.44 99.55 99.80 99.76 99.85 99.91 99.71 98.85 99.93 99.53 99.97
2 95.44 93.18 98.82 89.20 95.96 98.99 95.07 98.78 97.40 97.49 99.67
3 81.76 93.83 91.11 97.25 93.50 96.68 85.18 80.64 91.25 100.00 97.67
4 99.99 98.70 99.71 98.42 99.04 99.55 98.26 99.71 99.13 99.46 99.81
5 75.45 83.74 94.04 83.80 92.42 95.96 60.39 84.02 94.19 96.39 97.31
6 98.66 98.88 99.89 97.69 99.34 99.77 99.49 98.76 99.76 99.33 99.86
7 99.96 99.99 99.97 99.98 99.99 99.99 99.99 99.94 99.99 99.85 99.99
8 98.41 96.66 96.33 89.45 97.19 96.46 96.67 95.40 97.95 86.82 96.77
9 91.08 94.77 93.95 73.44 96.16 86.51 92.89 89.54 88.93 87.28 96.24
OA (%) 98.58 98.48 99.28 97.47 99.02 99.19 97.86 98.54 98.99 98.71 99.61
AA (%) 93.36 95.37 97.07 92.11 97.05 97.09 91.96 93.96 96.20 96.24 98.59
$\kappa$ 0.9812 0.9801 0.9905 0.9668 0.9872 0.9893 0.9718 0.9807 0.9868 0.9830 0.9949
@@
Compared with other methods, our proposed MTMSD achieves the highest OA, AA, and $\kappa$ on four datasets. According to the results, the CNN-based supervised methods, 2-D CNN, 3-D CNN, and SSRN, obtain good performance owing to their ability to capture local spatial information. Besides, since transformers are capable of capturing sequential information, transformer-based supervised methods, SF, SSFTT, and GAHT, also achieve competitive performance. Unsupervised methods, 3DCAE, 3DAES, and UMSDFL are proposed to tackle the problem of limited samples by learning representative features without any labeled samples. Limited by the model architecture, the features learned in an unsupervised manner are not discriminative enough for HSI classification lacking high-level information. Therefore, the performance of 3DCAE and 3DAES is even lower than that of some explicit learning methods. SpetralDiff introduces the diffusion model to HSI classification and achieves competitive performance due to the advantages of diffusion features. However, the feature used in the method is from a single timestep and a single layer with the loss of some key information, which limits the performance. Our proposed MTMSD explores the value of multi-timestep multi-stage diffusion features through class & timestep-oriented multi-stage feature purification and selective timestep feature fusion, effectively modeling complex spectral-spatial relations due to the adaptive integration of contextual semantics and textual details. Thus, our MTMSD outperforms all the previous methods on four datasets: Indian Pines, PaviaU, Houston 2018, and Longkou. Notably, the classification performance of the Houston 2018 Dataset is largely improved compared with the previous SOTA method in terms of OA (98.29% versus 96.69%), AA (96.04% versus 92.98%), and $\kappa$ (0.9777 versus 0.9570), which especially demonstrate our effectiveness.
Figure 8: Classification maps obtained by different methods on the WHU-Hi-Longkou dataset. (a) Ground truth. (b) 2-D CNN (OA=98.58%). (c) 3-D CNN (OA=98.48%). (d) SSRN (OA=99.28%). (e) SF (OA=97.47%). (f) SSFTT (OA=99.02%). (g) GAHT (OA=99.19%). (h) 3DCAE (OA=97.86%). (i) 3DAES (OA=98.54%). (j) UMSDFL (OA=98.99%). (k) SpectralDiff (OA=98.71%). (l) MTMSD (OA=99.61%).
IV-C2 Classification Results with Different Proportions of Training Samples
The classification results with different proportions of training samples are shown in Fig. 9. It can be observed that the performance increases with the percentages of training samples. Our method outperforms the compared method consistently in terms of OA on four datasets. Especially, using only 2% of the training sample, our method achieves comparable results to other methods using 5% of the training sample on the Houston 2018 dataset.
IV-D Ablation Studies
In this section, we analyze the effect of the components in our method.
IV-D1 Ablation for Class & Timestep-oriented Multi-Stage Feature Purification
Table IX presents the effect of class & timestep-oriented multi-stage feature purification (CTMSFP) across the four datasets. The multi-stage features extracted from the diffusion model are not entirely aligned with the HSI classification task. Therefore, the CTMSFP is proposed to reduce the redundancy of the features and maintain efficiency. All experiments are conducted on an RTX 3090 GPU, with a consistent batch size of 64 used during training. As shown in Table IX, the results demonstrate that CTMSFP significantly reduces the number of parameters and the GPU memory consumption during training. This reduction is due to the channel-wise purification performed by CTMSFP before the features are input into the model, which substantially decreases the size of both the input tensor and the model structure. Furthermore, the average inference time is reduced by 25%, effectively enhancing the model’s computational efficiency, while still ensuring a slight performance improvement. We also compare the performance and efficiency of another diffusion-based HSI classification model, SpectralDiff, which utilizes single-timestep single-stage diffusion features. Our method performs independent linear transformations on features from different timesteps, allowing tailored transformations for diverse multi-timestep feature patterns, resulting in a larger number of parameters. However, our method with CTMSFP significantly leads to performance, inference speed, and GPU memory consumption.